# Mitigate Language Priors in Large Vision-Language Models by Cross-Images Contrastive Decoding

### v1

@@ Abstract
Language priors constitute one of the primary causes of hallucinations in Large Vision-Language Models (LVLMs), driving the models to generate linguistically plausible yet visually inconsistent content. The language priors in LVLMs originate from the linguistic knowledge inherited from their pre-trained Large Language Model (LLM) backbone. Consequently, this characteristic is an intrinsic property of the model that remains independent of visual inputs. Inspired by the finding that language priors are consistent across images, we propose Cross-Image Contrastive Decoding (CICD), a simple yet effective training-free method to alleviate language priors in LVLMs. CICD first identifies essential and detrimental priors, and then employs contrastive decoding to eliminate the detrimental ones. This approach simultaneously prevents LVLMs from generating hallucinated content while maintaining textual fluency and coherence. Furthermore, the limited information overlap between images helps prevent visual information loss during contrastive decoding. We validate the effectiveness of CICD on four benchmarks with six LVLMs. Our experiments demonstrate that CICD performs remarkably well in mitigating language priors, especially in the image captioning task, where such priors are most pronounced. Code will be released once accepted.
Index Terms:
Large Vision-Language Model, Multi-Model Hallucination, Language Priors
I Introduction
In recent years, the remarkable success of Large Language Models (LLMs) [1; 2; 3; 4] has provided a viable technological pathway toward Artificial General Intelligence (AGI). Numerous studies [5; 6; 7; 8] have extended these models by incorporating visual inputs, resulting in Large Vision-Language Models (LVLMs) capable of visual comprehension. These models perform various tasks using information from visual inputs, such as image captioning [9; 10; 11; 12], visual question answering [13; 14; 15], and visual grounding [16; 17; 18]. Despite their fluency, LVLMs are prone to hallucinations [19; 20], generating text that contradicts visual evidence.
Hallucinations arise from various factors, most directly from language priors [12; 21; 22; 23], which cause LVLMs to struggle with accurately expressing their understanding of visual information. Since LVLMs are built upon LLMs, they inherently inherit the linguistic knowledge embedded in the LLMs, resulting in responses that reflect both visual cues and language priors. For example, when describing an office scene, an LVLM may incorrectly generate the word “computer” even if no such object is present in the image.
There are many approaches to alleviating language priors. Some works [24; 25; 12] construct specialized training datasets to reduce models’ reliance on language priors. However, the rapid iteration of model backbones limits the practical applicability of such training-based methods. Recently, a series of studies [21; 22; 23; 26] have used the contrastive decoding algorithm [27] to mitigate language priors, achieving remarkable effectiveness. Specifically, these works construct modified contexts by either perturbing [21; 22] or enhancing [23; 26] the visual features from the original context, and then apply contrastive decoding between the original and modified contexts (i.e., subtracting one logit distribution from the other) to reduce language priors.
These methods are sufficiently simple, training-free, and exhibit excellent generalizability. However, since the modified context is derived from the original context, it retains partial visual information, which can be degraded during contrastive decoding. Moreover, the perturbations introduced may corrupt the model’s features, leading to information distortion.
@@ I Introduction
Figure 1: A case of language prior. Adding the same prefix when captioning another image will leads to the same language priors.
The language priors are derived from LLM backbones and are unrelated to the input image. In other words, the language priors remain consistent across different images, as shown in Fig.1. Inspired by this observation, we propose Cross-Images Contrastive Decoding (CICD), a simple yet highly effective method for mitigating language priors in LVLMs. Concretely, we replace the input image with a different image to create a new context that shares the same language priors. We then subtract this logit distribution from the original one to alleviate the influence of language priors, while preserving the original visual information. Since different images contain limited overlapping visual content, our method effectively prevents visual information loss.
@@
Figure 2: Essential priors and detrimental priors. We random select 500 images from the MSCOCO dataset and visualize LLaVA-1.5’s output by t-SNE algorithm [28]. w/ Image represents regular outputs. w/o Image represents the outputs after removing input image. w/ Neg. Image represents replace the input image by the most dissimilar image in the dataset retrieved by CLIP [29].
Figure 3: A case of JS divergence between the logit distribution of two images.
@@
While language priors contribute to hallucinated content in LVLMs, they also play a role in enhancing output fluency. Therefore, we categorize language priors into two types: (1) essential priors, which are necessary for maintaining text fluency, and (2) detrimental priors, which introduce vision-irrelevant information. As shown in Fig. 2, essential priors demonstrate high consistency across different images, whereas detrimental priors exhibit substantial divergence when generating image-related content. Based on this phenomenon, we distinguish between essential and detrimental priors using cross-image consistency, as illustrated in Fig.3. We observe a distinct divergence in cross-image consistency between essential and detrimental priors. We compute the Jensen-Shannon (JS) divergence between the two logit distributions, and set a threshold to identify essential priors that should be retained.
@@
We validate the effectiveness of our method, CICD, on four benchmarks: CHAIR [10], POPE [14], AMBER [30], and DetailCaps [11]. To evaluate the generalizability of CICD, we conduct experiments on six LVLMs: InstructBLIP [6], LLaVA-1.5 [31], LLaVA-Next [32], Qwen-VL-Chat [7], Qwen2-VL [33], and Qwen2.5-VL [8]. Experimental results demonstrate that CICD effectively mitigates language priors in LVLMs, thereby reducing hallucinations. This improvement is particularly notable in the image captioning task, which typically suffers from stronger language priors due to its longer textual context.
@@
  • We conduct a thorough investigation of language priors and demonstrate that they are intrinsic, model-derived properties independent of visual inputs. Furthermore, we show that consistency across different images can be used to distinguish essential priors from detrimental ones.
  • We propose CICD (Cross-Images Contrastive Decoding), which leverages both the consistency of language priors across images and the divergence of inter-image visual features to mitigate linguistic priors while preserving visual information.
@@
  • Extensive experiments across four benchmarks and six model backbones demonstrate the superiority of our method in alleviating language priors, particularly in the image captioning task.
II Related Work
II-A Large Vision-Language Models
Typically, Large Vision-Language Models (LVLMs) [6; 31; 7; 34] consist of a visual encoder [29; 35] and a Large Language Model (LLM) backbone [2; 3]. The visual encoder extracts visual features, which are then aligned with text tokens through a fusion module to form the input context for the LLM. Through multi-stage training, including modality alignment and instruction tuning [5], LVLMs acquire the ability to perform various multimodal tasks such as image captioning [11], visual question answering [15], and visual grounding [18]. Despite significant advancements, LVLMs still suffer from severe hallucination problems, often generating content that contradicts the visual input.
II-B Hallucination in LVLMs
Originally, hallucinations [36; 37] were defined as undesirable outputs or non-factual information. In the multimodal domain, the study of hallucinations primarily focuses on cross-modal consistency [19; 20], requiring that LVLM-generated content remains faithful to the visual inputs. Multiple factors contribute to hallucinations in LVLMs, including biased visual attention [38; 39], attention misallocation [40; 41; 42], and language priors [21; 23; 22]. Among these, language priors are the most direct cause, making it difficult for LVLMs to accurately express their understanding of visual information.
II-C Language Priors
As LVLMs are built upon LLMs, they inherently inherit the linguistic knowledge embedded in LLMs, which are referred to as language priors [24; 25]. Language priors specifically manifest when LVLMs generate content that is linguistically plausible but inconsistent with visual evidence. Some works [12; 43; 44] meticulously design specialized training data to mitigate this issue. However, such training-based approaches exhibit limited generalizability and cannot be easily transferred to broader application scenarios. Recently, a series of training-free approaches have directly reduced language priors using contrastive decoding [27; 45], achieving remarkable performance. These methods construct an alternative logit distribution on top of the original one through techniques such as masking the image [21], perturbing the instruction [22], augmenting the vision input [23], or performing cross-modal conversion [26]. During decoding, the two logit distributions are contrasted to eliminate language priors.
III Preliminary and Motivation
LVLMs retain the core architecture of LLMs while augmenting it with a vision encoder for visual feature extraction and contextual integration. The input image is encoded into visual tokens and inserted into the textual context. After forward propagation, LVLMs project the outputs into a logit distribution to compute conditional probabilities and generate responses following an autoregressive paradigm:
$y_{t}\sim p_{\theta}(y_{t}\mid v,x,y_{<t})=\mathrm{softmax}\left[\mathrm{logit}_{\theta}(y_{t}\mid v,x,y{<t})\right]$ (1)
where $v$ and $x$ denote the visual input and textual input, $y_{t}$ denotes the token generated at the $t$-th step, and $y_{<t}=[y_{1},y_{2},\cdots,y_{t-1}]$ is the sequence of tokens generated before $y_{t}$. $\mathrm{logit}_{\theta}$ represents the logit distribution formulated by the LVLM with parameters $\theta$, and is normalized into conditional probabilities $p_{\theta}$.
Since LVLMs are fine-tuned from pre-trained LLMs to enhance fluent generation capabilities, they inherently retain prior knowledge that is independent of visual inputs, i.e., language priors. Concretely, tokens ranked higher in the LVLM’s distribution $\mathrm{logit}_{\theta}$ may be semantically plausible but objectively inconsistent with visual evidence. This phenomenon results in hallucinations, where the generated outputs are linguistically fluent yet visually contradictory.
Language priors can be directly suppressed by downweighting the corresponding tokens in the logit distribution. However, since these tokens cannot be explicitly identified, contrastive decoding [27; 21] has emerged as a widely adopted solution to mitigate language priors. This method constructs an auxiliary distribution with enhanced language priors, and then computes a new distribution by contrasting the original and auxiliary distributions:
$\mathrm{logit}_{\mathrm{CD}}(y_{t}\mid v,x,y_{<t})=$ $(1+\alpha)\,\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})$ (2)
$-\alpha\,\mathrm{logit}_{C}$
@@ III Preliminary and Motivation
where $\mathrm{logit}_{C}$ is the auxiliary distribution, and $\mathrm{logit}_{\mathrm{CD}}$ is the final distribution used for next-token prediction.
The core principle of contrastive decoding lies in constructing a logit distribution that preserves consistent language priors while excluding consistent visual information. This ensures that visual information is not inadvertently removed along with language priors during contrastive decoding. However, since LVLMs are fine-tuned using multimodal data, removing the visual input distorts the resulting logit distribution, leading to inconsistent language priors (corresponding experiments are presented in Sec.V-G).
We demonstrate that language priors are inherent characteristics of LVLMs and remain consistent across different images, as shown in Fig.1. An intuitive approach is to use other images to construct a logit distribution that shares the same language priors but contains different visual information. Accordingly, Eq.2 is modified as:
$\mathrm{logit}_{\mathrm{CD}}(y_{t}\mid v,v^{\prime},x,y_{<t})=$ $(1+\alpha)\,\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})$ (3)
$-\alpha\,\mathrm{logit}_{\theta}(y_{t}\mid v^{\prime},x,y_{<t})$
where $v\textquoteright$ ’ is a different image from $v$. $\mathrm{logit}_{\theta}(y_{t}\mid v,x,y{<t})$ encodes both the visual information $V$ from image $v$ and the language priors $T$, which we denote as $V\circ T$. The effectiveness of contrastive decoding can be formulated as:
[…]
where $V\textquoteright$ ’ denotes the visual information of image $v\textquoteright$ ’, and $I(V;V\textquoteright)$ ’ ) denotes the consistency between $V$ and $V\textquoteright$ ’. This approach achieves two objectives simultaneously: (1) precise elimination of language priors, and (2) full preservation of visual information integrity.
IV Method
@@
Figure 4: The input image and another distinct image are fed into LVLM respectively with identical text-prefix. We calculate the JS discrepancy between the two logit distributions to discriminate essential priors and detrimental priors. For detrimental priors, we perform contrastive decoding to alleviate language priors, while applying regular decoding for essential priors. The final output of this case is shown in the right part.
IV-A Image Retrieval
To prevent visual information loss during contrastive decoding, we select an image that has limited information overlap with the input image. Since visual features are extracted by the LVLM’s vision encoder, an effective strategy is to measure the similarity between two images using the vision encoder to estimate their information overlap. In other words, the image with the lowest similarity to the input image is selected as the reference image. For better generalizability, we adopt the pre-trained CLIP [29] as the retriever. We vectorize all images in the dataset in advance, and during inference, retrieve the one with the minimum cosine similarity to the input image.
Although vector retrieval is time-efficient given the computational cost of LVLMs, practical deployment still requires loading an additional retriever and maintaining a vector database. As illustrated in Fig.1, we observe that language priors primarily manifest in content words such as objects, attributes, and relations. These visual elements tend to exhibit minimal information overlap across natural images. As a result, randomly selecting an image can be an effective and lightweight alternative in practical implementations. We further discuss this in Sec.V-G.
To ensure reproducibility, we adopt the retrieval-based approach for deterministic image selection throughout our experiments.
IV-B Identifying Detrimental Priors
Although language priors can lead LVLMs to generate hallucinated content, they are still essential for producing fluent and coherent text. Therefore, we categorize language priors into two types: essential priors and detrimental priors, where only the latter need to be eliminated. We define essential priors as purely linguistic attributes that are entirely independent of visual information. Specifically, for essential priors, the logit distributions produced by LVLMs should remain highly consistent across diverse visual inputs. This intrinsic property enables a clear distinction between essential and detrimental priors.

### v2

@@ Abstract
Language priors are a major cause of hallucinations in Large Vision-Language Models (LVLMs), often leading to text that is linguistically plausible but visually inconsistent. Recent work explores contrastive decoding as a training-free solution, but these methods typically construct negative visual contexts from the original image, resulting in visual information loss and distorted distribution. Motivated by the observation that language priors stem from the LLM backbone and remain consistent across images, we propose Cross-Images Contrastive Decoding (CICD), a simple yet effective training-free method that uses different images to construct negative visual contexts. We further analyze the cross-image behavior of language priors and introduce a distinction between essential priors (supporting fluency) and detrimental priors (causing hallucinations), enabling selective suppression. By selectively preserving essential priors and suppressing detrimental ones, our method reduces hallucinations while maintaining coherent and fluent language generation. Experiments on four benchmarks and six LVLMs across three model families confirm the effectiveness and generalizability of CICD, especially in image captioning, where language priors are particularly pronounced. Code will be released upon acceptance.
Mitigate Language Priors in Large Vision-Language Models by Cross-Images Contrastive Decoding
Jianfei Zhao1,2, Feng Zhang1, Xin Sun1, Chong Feng1,3,thanks: Corresponding Autor: chongfeng@bit.edu.cn , 1School of Computer Science and Technology, Beijing Institute of Technology, Beijing, China, 2Zhongguancun Academy, Beijing, China, 3Southeast Academy of Information Technology, Beijing Institute of Technology, Fujian, China,
1 Introduction
@@
Figure 1: A case of language prior. Adding the same prefix when captioning another image will leads to the same language priors. The model used is LLaVA-1.5.
Recent advances in Large Language Models (LLMs) Brown et al. (2020); Touvron et al. (2023); Chiang et al. (2023); Bai et al. (2023a) have opened promising pathways toward Artificial General Intelligence (AGI). Building on these developments, numerous studies Liu et al. (2023); Dai et al. (2023); Bai et al. (2023b, 2025) have extended LLMs by integrating visual inputs, giving rise to Large Vision-Language Models (LVLMs) capable of understanding visual content. They are capable of addressing a broad spectrum of vision-language tasks, such as image captioning Jing et al. (2024); Rohrbach et al. (2018); Ye et al. (2025); Chen et al. (2025), visual question answering (VQA) Lovenia et al. (2024); Li et al. (2023b); Yu et al. (2024), and visual grounding Rasheed et al. (2024); You et al. (2024); Lu et al. (2024).
Despite impressive capabilities, LVLMs remain susceptible to hallucinations Li et al. (2023c); Wang et al. (2023), often generating text that contradicts the visual input. These hallucinations stem from various sources, with language priors Chen et al. (2025); Leng et al. (2024); Wang et al. (2024c); Zhu et al. (2024b) being the most fundamental underlying cause. Since LVLMs are built upon LLMs, they inherently inherit the linguistic knowledge encoded in the LLMs. As a result, their outputs are influenced not only by visual evidence but also by pre-learned language patterns. This often leads to a tendency to rely on memorized linguistic associations when responding to user queries. For instance, when describing an office scene, an LVLM may erroneously mention a "computer" even if no such object is present in the image.
@@
Figure 2: Essential priors (left) and detrimental priors (right). We randomly select 500 images from the MSCOCO dataset and visualize LLaVA-1.5’s output by the t-SNE algorithm van der Maaten and Hinton (2008). w/ Image represents regular outputs. w/o Image represents the outputs after removing the input image. w/ Neg. Image represents replacing the input image by the most dissimilar image in the dataset retrieved by CLIP Radford et al. (2021).
Methods for mitigating language priors in LVLMs can generally be classified into two categories. The first involves constructing specialized training datasets to reduce the model’s dependence on linguistic biases Zhibo et al. (2023); Wu et al. (2022); Chen et al. (2025), though such approaches often incur substantial training costs. In this work, we focus on plug-and-play, training-free methods. Many recent studies adopt contrastive decoding Li et al. (2023a), which constructs negative contexts to expose and suppress language priors. Some approaches directly perturb the input image Leng et al. (2024); Wang et al. (2024c) to alter the visual context, while others enhance the visual features extracted from the original image Zhu et al. (2024b); Zhang et al. (2025). These methods typically compute the difference between the logit distributions obtained from the original and modified (i.e., negative) contexts to reduce the model’s reliance on language priors. Following the paradigm introduced in VCD Leng et al. (2024), these methods routinely modify the original image to generate negative visual contexts. However, this prevailing design choice raises a critical question: Is constructing negative contexts through perturbations of the original image truly the most effective strategy for inducing and mitigating language priors?
We argue that the answer is unequivocally negative. Constructing negative contexts based on the original image can lead to two main issues: (1) Visual information loss, where the negative visual context retains partial visual information, which may be degraded during contrastive decoding. (2) Distorted distribution, where perturbing the model’s input to induce priors may disrupt its internal representations, ultimately resulting in a distorted logit distribution. This analysis raises an important question: How can the above issues be avoided when constructing negative contexts?
In fact, simply taking another image to construct negative context is sufficient. This insight is based on the thought that the language priors guiding step-wise token generation in LVLMs originate from the LLM backbone and are largely independent of the input image. As shown in Fig.1, LVLM produces highly consistent outputs when given completely different images but the same prefix, which strongly supports our hypothesis. Building on this, we propose Cross-Images Contrastive Decoding (CICD), a simple yet effective method for mitigating language priors in LVLMs. Instead of modifying the original image, CICD constructs the negative context using a different image that shares the same language priors. Our method avoids the two key issues present in previous approaches. Since the overlap between the original and the new image is limited, our approach avoids introducing critical visual information into the negative context. Moreover, it enables the model to reveal intrinsic priors, maintaining the model’s natural inference process without injecting external bias. The formal analysis of our method is presented in Sec. 3.2.
Figure 3: A case of JS divergence between the logit distributions of two images.
@@
Focusing on language priors, we further explore whether all such priors should be strictly eliminated. We argue that the answer is negative. Language priors can be divided into two categories: (1) essential priors, which are necessary for maintaining textual fluency, and (2) detrimental priors, which introduce information unrelated to the visual input. Clearly, only the detrimental priors should be removed. This raises a critical question: How can we accurately distinguish between the two and selectively eliminate only the detrimental ones? As shown in Fig.2, essential priors exhibit high consistency across different images, whereas detrimental priors show substantial divergence when generating image-related content. Based on this observation, we propose to distinguish essential from detrimental priors using cross-image consistency, as illustrated in Fig.3. Specifically, we compute the Jensen–Shannon (JS) divergence between the logit distributions of different visual contexts, and apply a stable threshold to identify essential priors that should be retained.
@@
We validate the effectiveness of CICD on four widely used benchmarks: CHAIR Rohrbach et al. (2018), POPE Li et al. (2023b), AMBER Wang et al. (2024a), and DetailCaps Ye et al. (2025). To assess the generalizability of our approach, we conduct experiments across six LVLMs from three different model families: InstructBLIP Dai et al. (2023), LLaVA-1.5 Liu et al. (2024c), LLaVA-Next Liu et al. (2024d), Qwen-VL-Chat Bai et al. (2023b), Qwen2-VL-Instruct Wang et al. (2024b), and Qwen2.5-VL-Instruct Bai et al. (2025). Experimental results show that CICD consistently reduces hallucinations by effectively mitigating language priors in LVLMs, outperforming state-of-the-art methods across all benchmarks and models. The improvement is particularly pronounced in the image captioning task, which tends to suffer more from language priors due to its longer textual outputs.
@@
  • We conduct an in-depth analysis of existing contrastive decoding methods and identify their inherent limitations rooted in the reliance on the original image. We further verify the consistency of language priors across images and propose a stable approach for distinguishing essential priors from detrimental ones.
  • Based on this analysis, we propose CICD (Cross-Images Contrastive Decoding), which leverages retrieved images to induce language priors while selectively preserving essential priors during the mitigation process.
@@
  • Extensive experiments on multiple benchmarks and LVLMs demonstrate the superiority of our method in alleviating language priors, particularly in the image captioning task.
2 Related Work
Due to space limitations, we provide a brief overview here. A more detailed discussion of related work can be found in Sec.A.
2.1 Hallucination in LVLMs
Hallucinations in LVLMs primarily concern cross-modal consistency Li et al. (2023c); Wang et al. (2023), which requires that the generated content remains faithful to the visual input. Among the various causes of hallucination, language priors Zhibo et al. (2023); Wu et al. (2022) are among the most fundamental. These priors originate from the LLM’s pretraining corpus and guide generation based on learned language patterns, which can produce fluent yet visually inconsistent outputs. Several approaches Chen et al. (2025); Liu et al. (2024a); Gunjal et al. (2024) attempt to mitigate language priors by constructing specialized training datasets, but these methods often incur significant additional training costs. Alternatively, recent training-free methods have explored contrastive decoding Li et al. (2023a); Chuang et al. (2024), typically constructing negative visual contexts based on the original image Leng et al. (2024); Wang et al. (2024c); Zhu et al. (2024b); Zhang et al. (2025) in an attempt to elicit and suppress language priors.
In this work, we identify critical limitations of these approaches that stem from their reliance on modifying the original image. Specifically, such modifications may lead to visual information loss, biased language priors, and distribution mismatch. We further conduct an in-depth analysis of the cross-image consistency of language priors and, based on this insight, propose CICD. Additionally, we introduce a method for distinguishing between essential priors and detrimental priors, enabling fine-grained and precise mitigation of language priors in LVLMs.
3 Preliminary
Figure 4: We input the target image and a distinct image into the LVLM separately. During the step-by-step generation of the response, we compute the Jensen-Shannon (JS) divergence between the two logit distributions at each step to distinguish essential priors from detrimental ones. For detrimental priors, we apply contrastive decoding to suppress harmful language biases, while regular decoding is used for essential priors.
3.1 Contrastive Decoding
LVLMs take as input a concatenation of vision tokens from a vision encoder and an encoded text prompt. After forward propagation, the model projects the output into a logit distribution to compute conditional probabilities and generate responses in an autoregressive manner:
$y_{t}\sim p_{\theta}(y_{t}\mid v,x,y_{<t})$ (1)
$=$ $\ \mathrm{softmax}\left[\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})\right]$
Here, $v$ and $x$ represent the visual and textual inputs, respectively. $\mathrm{logit}_{\theta}$ represents the logit distribution formulated by the LVLM with parameters $\theta$.
Since LVLMs adopt pretrained LLMs as their backbone, they inherently inherit language priors that are independent of the visual input. As a result, tokens ranked higher in the output distribution $\mathrm{logit}_{\theta}$ may appear semantically plausible but can be objectively inconsistent with the visual evidence. This constitutes a fundamental cause of hallucinations. A commonly used approach to mitigate such priors is contrastive decoding Li et al. (2023a); Leng et al. (2024), which modifies the original image to construct a native visual context that elicits language priors. These priors are then suppressed by subtracting the logit distribution derived from the native context from that of the original context.
$\mathrm{logit}_{\mathrm{CD}}(y_{t}\mid v,x,y_{<t})$ (2)
$=$ $(1+\alpha)\,\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})-\alpha\,\mathrm{logit}_{N}$
@@ 3.1 Contrastive Decoding
Here, $\mathrm{logit}_{N}$ denotes the logit distribution obtained from the native context, and $\mathrm{logit}_{\mathrm{CD}}$ represents the final distribution used for next-token prediction.
3.2 Formal Analysis
The core component of contrastive decoding lies in the construction of the native context. Our preliminary investigation reveals that existing methods, which modify the original image, introduce several limitations (see Sec.1). Further observations and analysis show that language priors are inherent characteristics of LVLMs and remain consistent across different images, as illustrated in Fig.1. This insight leads to a natural alternative, where the native context is constructed using an image distinct from the original. Under this design, Eq. 2 is refined as:
$\mathrm{logit}_{\mathrm{CD}}(y_{t}\mid v,x,y_{<t})$ (3)
$=$ $\ (1+\alpha)\,\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})$
$\quad-\alpha\,\mathrm{logit}_{\theta}(y_{t}\mid v^{\prime},x,y_{<t})$
where $v\textquoteright$ ’ is a different image from $v$. We further provide a formal analysis of the effectiveness of this approach.
Most existing methods employ sampling decoding, where a token is randomly selected according to the logit distribution. We define the set of plausible tokens under a given logit distribution as $S=V\cup T$, where $V$ represents tokens related to visual content, and $T$ represents tokens unrelated to the image, i.e., those driven by language priors. The negative context constructed by a different image yields a new logit distribution, whose plausible token set is $S\textquoteright=V\textquoteright\cup T$, given that language priors remain unchanged across images. The contrastive decoding operation can thus be interpreted as subtracting one plausible token set from another:
$S-S^{\prime}=(V\cup T)-(V^{\prime}\cup T)=V\mbox{ if }V^{\prime}\cap V=\emptyset$ (4)
In other words, this subtraction removes all language priors that are unrelated to the visual input. This approach achieves three objectives simultaneously: (1) retention of all essential visual information, (2) precise elimination of language priors, and (3) compatibility with the input distribution expected by LVLMs.
4 Method
@@
4.1 Image Retrieval
As discussed earlier, it is essential to select an image with limited information that overlaps with the input image. Since visual features are extracted by the vision encoder, a natural strategy is to measure the similarity between two images using the encoder itself to estimate their overlap. In other words, the image with the lowest similarity to the input image is selected as the reference. To ensure better generalizability, we adopt the pretrained CLIP model Radford et al. (2021) as the retriever. All images in the dataset are vectorized in advance, and during inference, the image with the lowest cosine similarity to the input is retrieved as the contrastive image.
Although vector retrieval is relatively efficient given the computational cost of LVLMs, practical deployment still requires loading an additional retriever and maintaining a vector database. As illustrated in Fig.1, we observe that language priors primarily manifest in content words such as objects, attributes, and relationships. These visual elements tend to exhibit minimal overlap across natural images. Therefore, randomly selecting an image can serve as an effective and lightweight alternative for practical applications. We provide further discussion in Sec.C.2. Note that, to ensure reproducibility, we adopt the retrieval-based strategy for consistent and deterministic image selection throughout all experiments.
4.2 Identifying Detrimental Priors
Although language priors may cause LVLMs to generate hallucinated content, they remain essential for producing fluent and coherent text. To address this, we categorize language priors into two types: essential priors and detrimental priors, with only the latter requiring elimination. Essential priors are defined as purely linguistic patterns that are entirely independent of visual input. As discussed in Sec. 1, the logit distributions associated with essential priors should remain highly consistent across diverse visual inputs. This intrinsic property provides a reliable basis for distinguishing between essential and detrimental priors.